OpenStudy (anonymous):

how do you solve 20*2^(13/s)

OpenStudy (anonymous):

what? @Danielayolo

OpenStudy (tkhunny):

You don't. It's not an equation. Should there be "=" in there, somewhere?

OpenStudy (anonymous):

NORHING\

OpenStudy (anonymous):

20*2^(13/s)=y

OpenStudy (anonymous):

@tkhunny @Danielayolo @reemii

OpenStudy (reemii):

you want to find s as a function of y ?

OpenStudy (reemii):

do "as usual", at some point you will need to take the logarithm in base 2 on both sides.

OpenStudy (tkhunny):

\(20\cdot 2^{13/s} = y\) Divide by 20 \(2^{13/s} = \dfrac{y}{20}\) Now what? I recommend a logarithm.

OpenStudy (anonymous):

@tkhunny like s=ln(8192)/ln(y)-ln(20)

OpenStudy (tkhunny):

No. I have no idea how you managed that. Please provide intermediate steps. \(\log\left(2^{13/s}\right) = \log(y/20)\) \(\dfrac{13}{s}\log(2) = \log(y/20)\) Now what?

OpenStudy (anonymous):

@tkhunny remove the log?

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

remove \(log_2(2)\)? . it is equal to 1 so i think what you mean is correct. then it's just a matter of isolating \(s\) and obtaining an equation of the form \(s=....\). easy now.

OpenStudy (anonymous):

\[13/s=\] idk how to do the other side :(

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

does not change.

OpenStudy (anonymous):

so its \[13/s=\log(y/20)\]

OpenStudy (reemii):

yes, now you have \(s=13/\log(y/20)\). you're done.

OpenStudy (anonymous):

@reemii how do i get an exact answer?

OpenStudy (reemii):

if yuo don't know the value of \(y\), there's nothing else to do here.

OpenStudy (tkhunny):

I used Base 10 logs. No removing ever should occur.